# Cavity-induced coherent magnetization and polaritons in altermagnets

### v1

@@ Abstract
Introduction.—Altermagnets represent a recently identified class of magnetic materials characterized by collinear spin arrangements and unorthodox spin symmetry rules [1, 2, 3]. Unlike ferromagnets or antiferromagnets, they exhibit neither net magnetization nor conventional Néel order [1, 3, 2, 4, 5, 6, 7, 8, 9, 10, 11], but remain invariant under combined spin and specific crystal rotations [1]. This symmetry breaking yields nonrelativistic spin splitting, highly anisotropic spin transport, and distinctive magnetic responses despite vanishing total moment. These properties make altermagnets promising platforms for next-generation spintronic technologies [4, 5, 7, 8, 9, 12].
Their zero-moment ground state minimizes stray fields—desirable in low-noise and interference-sensitive applications—but hinders magnetic control through external fields [1, 3, 2, 13, 6]. Static methods to induce magnetization have been demonstrated [14, 15, 16, 17, 18], yet dynamical generation remains open [19, 20, 21]. Dynamical routes offer access to far-from-equilibrium regimes, enabling ultrafast and reversible spin control without altering intrinsic material properties.
Ultrafast optical and cavity quantum electrodynamics techniques provide powerful tools for manipulating quantum materials [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. In altermagnets, optical driving can reorient spins, tune orbitals, and reshape bands [36, 37, 38, 39]. Strong light-matter coupling in optical cavities creates hybrid polaritonic states that reshape electronic, vibrational, or magnetic excitations [23, 40, 41, 42, 40, 41, 42, 43, 32, 30, 44, 45, 46, 47, 48, 49]. In magnetic systems, cavity photons can mediate long-range interactions and tune collective modes, as recently demonstrated for magnon-polaritons in altermagnets [50]. However, direct control of altermagnetic properties using driven cavities has not yet been addressed.
@@
Figure 1: (Left) Spin-resolved band structure of a $d_{x^{2}-y^{2}}$-wave altermagnet along a high-symmetry $\vec{k}$-path, showing momentum-dependent spin splitting between spin-up and spin-down bands. The inset depicts the collinear antiparallel spin arrangement characteristic of altermagnetic order. (Right) Schematic of the proposal: a $d_{x^{2}-y^{2}}$-wave altermagnet embedded in a coherently driven optical cavity. The coherent laser drive excites cavity photons, which couple to staggered sublattices and enable cavity-induced coherent magnetization.
Model.—We consider a 2D $d$-wave altermagnet described by a two-sublattice tight-binding model on a square lattice, where $\uparrow,\downarrow$ label spin degrees of freedom arising from sublattice distinctions. The model incorporates direction-dependent hopping (staggered fields), electron-photon interactions, and a driven cavity photon mode; see Fig. 1(right). In the absence of the cavity, the model captures the key features of $d$-wave altermagnetism. Candidate materials include KV2Se2O, RbV2Te2O, $\kappa$-Cl, RuO2, and FeSe [51, 52, 53, 12, 4, 9, 54, 1, 55, 56], several of which have been experimentally confirmed to exhibit altermagnetic order and can be exfoliated to mimic the 2D geometry used here. The total Hamiltonian reads as $H=H_{\rm e}+H_{\rm ph}+H_{\rm e\text{-}ph}$, where $H_{\rm e}$ describes electrons, $H_{\rm ph}$ photons, and $H_{\rm e\text{-}ph}$ their coupling.
@@
and dispersion $\mathcal{E}_{\vec{k}}=\epsilon_{\vec{k}}\pm\Delta_{\vec{k}}$; see Fig. 1(left)  [^1] In our framework, we introduce the notion of spin magnetization using the Pauli matrix $\sigma_{z}$ in orbital space, assigning $\sigma_{z}=+1$ to the first orbital and $\sigma_{z}=-1$ to the second. The total pseudospin magnetization per unit cell is defined as $n_{\uparrow}-n_{\downarrow}\propto\int_{\mathrm{BZ}}\frac{d^{2}k}{(2\pi)^{2}}\langle\sigma_{z}\rangle_{\vec{k}}$. Due to the antisymmetry of the spin-splitting term, $\Delta_{(k_{x},k_{y})}=-\Delta_{(k_{y},k_{x})}$, under interchange of $k_{x}$ and $k_{y}$, and the corresponding symmetry of the Brillouin zone, the integral evaluates to zero. Consequently, in the absence of cavity photons, the momentum-dependent spin splitting does not produce a net spin magnetization, implying that spin densities remain perfectly balanced across the Brillouin zone, as shown in Fig. 1 (left).. The antisymmetry $\Delta_{(k_{x},k_{y})}=-\Delta_{(k_{y},k_{x})}$ enforces zero net magnetization in equilibrium [12, 4, 9, 54, 1, 55, 56, 58]. In the following, we demonstrate that this antisymmetry is disrupted once cavity photons are coherently driven and coupled to the spin-polarized electrons. The resulting photon-assisted processes generate a momentum-space imbalance in spin densities, thereby inducing a finite magnetization in the altermagnet.
@@
with dimensionless couplings $\lambda_{\rm linear}=g_{\rm linear}/\Delta_{0}$ and $\lambda_{\rm quadratic}=g_{\rm quadratic}/\Delta_{0}$ set by bond geometry and polarization [47, 46, 34, 61, 22]. This local interaction term selectively couples photon displacement operators to the on-site electron density, while conserving spin. The staggered potential $\Delta_{0}$ breaks local inversion symmetry between sublattices, making the coupling spin selective. This photon-induced imbalance between spin sectors drives a finite pseudospin magnetization, as illustrated schematically in Fig. 1. In a conventional collinear antiferromagnet ($\Delta_{0}=0$), no such asymmetry exists, and cavity driving cannot break local inversion symmetry between sublattices on the same square lattice. In the thermodynamic limit, quantum fluctuations of the cavity field scale as $1/\sqrt{N}$ and can be neglected, enabling a mean-field treatment of $H_{\rm e\text{-}ph}$.
@@
We focus on the strong altermagnetic limit $\Delta_{0}\approx J$, where the spin splitting is comparable to the bandwidth, allowing analytical insight. For weaker $\Delta_{0}$, qualitative behavior remains similar but requires numerical evaluation. Independent of the $\Delta_{0}>0$ value, the net magnetization remains zero in the absence of photons. In the strong altermagnetic limit $\Delta_{0}\sim J$, an analytical expression for the NESS spin densities ($\dot{n}_{\sigma}=0$) is obtained at $T\to 0$: $\frac{1}{N}\sum_{\vec{k}}N_{\vec{k}}^{\sigma}\approx\frac{1}{\pi}\arccos\left(\frac{\sigma\,\widetilde{g}\,q}{4J}\right)$, which shows how the cavity-induced photon displacement $q$ directly sets the fraction of occupied states for each spin.
Figure 2: Time evolution of the spin imbalance $n_{\uparrow}(t)-n_{\downarrow}(t)$ for (a) varying $g_{\rm linear}$ with $g_{\rm quadratic}=0$ and (b) varying $g_{\rm quadratic}$ with $g_{\rm linear}=0$. All couplings reach a NESS, with stronger couplings accelerating the onset and enhancing the steady-state magnetization, reflecting cavity-induced sublattice symmetry breaking. Insets reveal coherent long-time oscillations for a few couplings, characteristic of the NESS. Parameters: $\omega_{\rm photon}/J=0.5$, $\omega_{\rm laser}=\omega_{\rm photon}$, $\mathcal{E}_{\rm laser}/J=0.1$, $\gamma_{\rm electron}=\gamma_{\rm photon}=\kappa=0.02J$.

### v2

@@ Abstract
Introduction.—Altermagnets represent a recently identified class of magnetic materials characterized by collinear spin arrangements and unorthodox spin symmetry rules [1, 2, 3]. Unlike ferromagnets or antiferromagnets, they exhibit neither net magnetization nor conventional Néel order [1, 3, 2, 4, 5, 6, 7, 8, 9, 10, 11], but remain invariant under combined spin and specific crystal rotations [1]. This symmetry breaking yields nonrelativistic spin splitting, highly anisotropic spin transport, and distinctive magnetic responses despite vanishing total moment. These properties make altermagnets promising platforms for next-generation spintronic technologies [4, 5, 7, 8, 9, 12, 13].
Their zero-moment ground state minimizes stray fields—desirable in low-noise and interference-sensitive applications—but hinders magnetic control through external fields [1, 3, 2, 14, 6]. Static methods to induce magnetization have been demonstrated [15, 16, 17, 18, 19], yet dynamical generation remains open [20, 21, 22, 23]. Dynamical routes offer access to far-from-equilibrium regimes, enabling ultrafast and reversible spin control without altering intrinsic material properties.
Figure 1: Schematic illustration of a laser-driven cavity setup hosting (left) a conventional $s$-wave antiferromagnet and (right) a $d_{x^{2}-y^{2}}$-wave altermagnet. The coherent laser drive excites cavity photons, which couple to spin textures via light-matter interaction. In the antiferromagnetic case, the spin populations remain nearly balanced under periodic driving, yielding minimal nonequilibrium polarization. In contrast, the anisotropic altermagnetic order due to staggered sublattices leads to a pronounced photoinduced imbalance between spin populations, as shown by the time evolution of the spin-resolved occupation difference $n_{\uparrow}(t)-n_{\downarrow}(t)$ (bottom panels).
Ultrafast optical and cavity quantum electrodynamics techniques provide powerful tools for manipulating quantum materials [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. In altermagnets, optical driving can reorient spins, tune orbitals, and reshape bands [42, 43, 44, 45]. Strong light-matter coupling in optical cavities creates hybrid polaritonic states that reshape electronic, vibrational, or magnetic excitations [25, 46, 47, 48, 46, 47, 48, 49, 34, 32, 50, 51, 52, 53, 54, 55]. In magnetic systems, cavity photons can mediate long-range interactions and tune collective modes, as recently demonstrated for magnon-polaritons in altermagnets [56]. However, direct control of altermagnetic properties using driven cavities has not yet been addressed.
@@
Model.—We consider a 2D $d$-wave altermagnet described by a two-sublattice tight-binding model on a square lattice, where $\uparrow,\downarrow$ label spin degrees of freedom arising from sublattice distinctions. The model incorporates direction-dependent hopping (staggered fields), electron-photon interactions, and a driven cavity photon mode; see Fig. 1. In the absence of the cavity, the model captures the key features of $d$-wave altermagnetism. Candidate materials include KV2Se2O, RbV2Te2O, $\kappa$-Cl, RuO2, and FeSe [57, 58, 59, 12, 4, 9, 60, 1, 61, 62], several of which have been experimentally confirmed to exhibit altermagnetic order and can be exfoliated to mimic the 2D geometry used here. The total Hamiltonian reads as $H=H_{\rm e}+H_{\rm ph}+H_{\rm e\text{-}ph}$, where $H_{\rm e}$ describes electrons, $H_{\rm ph}$ photons, and $H_{\rm e\text{-}ph}$ their coupling.
@@
and dispersion $\mathcal{E}_{\vec{k}}=\epsilon_{\vec{k}}\pm\Delta_{\vec{k}}$  [^1] In our framework, we introduce the notion of spin magnetization using the Pauli matrix $\sigma_{z}$ in orbital space, assigning $\sigma_{z}=+1$ to the first orbital and $\sigma_{z}=-1$ to the second. The total pseudospin magnetization per unit cell is defined as $n_{\uparrow}-n_{\downarrow}\propto\int_{\mathrm{BZ}}\frac{d^{2}k}{(2\pi)^{2}}\langle\sigma_{z}\rangle_{\vec{k}}$. Due to the antisymmetry of the spin-splitting term, $\Delta_{(k_{x},k_{y})}=-\Delta_{(k_{y},k_{x})}$, under interchange of $k_{x}$ and $k_{y}$, and the corresponding symmetry of the Brillouin zone, the integral evaluates to zero. Consequently, in the absence of cavity photons, the momentum-dependent spin splitting does not produce a net spin magnetization, implying that spin densities remain perfectly balanced across the Brillouin zone, as shown in Fig. 1 (left).. The antisymmetry $\Delta_{(k_{x},k_{y})}=-\Delta_{(k_{y},k_{x})}$ enforces zero net magnetization in equilibrium [12, 4, 9, 60, 1, 61, 62, 64]. In the following, we demonstrate that this antisymmetry is disrupted once cavity photons are coherently driven and coupled to the spin-polarized electrons. The resulting photon-assisted processes generate a momentum-space imbalance in spin densities, thereby inducing a finite magnetization in the altermagnet.
@@
with dimensionless couplings $\lambda_{\rm linear}=g_{\rm linear}/\Delta_{0}$ and $\lambda_{\rm quadratic}=g_{\rm quadratic}/\Delta_{0}$ set by bond geometry and polarization [53, 52, 36, 67, 24]. This local interaction term selectively couples photon displacement operators to the on-site electron density, while conserving spin. The staggered potential $\Delta_{0}$ breaks local inversion symmetry between sublattices, making the coupling spin selective. This photon-induced imbalance between spin sectors drives a finite pseudospin magnetization, as illustrated schematically in Fig. 1(right). In a conventional collinear antiferromagnet ($\Delta_{0}=0$) with the same lattice configuration, see Fig. 1(left), no such asymmetry exists, and cavity driving cannot break local inversion symmetry between sublattices on the same square lattice. In the thermodynamic limit, quantum fluctuations of the cavity field scale as $1/\sqrt{N}$ and can be neglected, enabling a mean-field treatment of $H_{\rm e\text{-}ph}$.
@@
We focus on the strong altermagnetic limit $\Delta_{0}\approx J$, where the spin splitting is comparable to the bandwidth, allowing analytical insight. For weaker $\Delta_{0}$, qualitative behavior remains similar but requires numerical evaluation. Independent of the $\Delta_{0}>0$ value, the net magnetization remains zero in the absence of photons. In the strong altermagnetic regime, where $\Delta_{0}\sim J$, an analytical expression for the NESS spin densities ($\dot{n}_{\sigma}(t)=0$) can be derived in the zero-temperature limit ($T\!\to\!0$):
$\frac{1}{N}\sum_{\vec{k}}N_{\vec{k}}^{\sigma}(t)\simeq\frac{1}{\pi}\arccos\!\left(\frac{\sigma\,\widetilde{g}(t)\,q(t)}{4J}\right).$ (8)
This relation explicitly shows that the cavity-induced photon displacement $q(t)$ directly controls the fraction of occupied electronic states for each spin species, thereby generating a spin imbalance and, consequently, a cavity-induced magnetization.
Figure 2: Time evolution of the spin imbalance $n_{\uparrow}(t)-n_{\downarrow}(t)$ for (a) varying $g_{\rm linear}$ with $g_{\rm quadratic}=0$ and (b) varying $g_{\rm quadratic}$ with $g_{\rm linear}=0$. All couplings reach a NESS, with stronger couplings accelerating the onset and enhancing the steady-state magnetization, reflecting cavity-induced sublattice symmetry breaking. Insets display coherent long-time oscillations, characteristic of the NESS, for several coupling strengths, obtained after subtracting their respective NESS values to align all traces around zero for comparison. Parameters: $\omega_{\rm photon}/J=0.5$, $\omega_{\rm laser}=\omega_{\rm photon}$, $\mathcal{E}_{\rm laser}/J=0.1$, $\gamma_{\rm electron}=\gamma_{\rm photon}=\kappa=0.02J$.
@@
Data availability.— The data that support the findings of this article are openly available at [83].